1. Filter module (FM1) on PRCL and MICH show significant delay while enabling and disabling.
2. I tried to fix PMC alignment (PMC trans was 0.76). I was not able to get PMC trans more than 0.79.
PMC has been this way since yesterday.
3. MICH is still bright when locked (ASDC_OUT reads 0.08 for dark and 2.0 for bright). We suspect it is because of the AS55_I error offset that persists even after running LSCoffsets script.
4. PRMI shows some dither at 3Hz when locked.
PMC is fixed with 0.84 in transmission. It was misaligned in pitch (fixing this increased PMC_trans to 0.822 from 0.773) and Koji also touched the wave plate and polarizer (changed PMC_trans to 0.845).